# Variational inequalities (VIs) have attracted researchers’ attention in various fields for more than half a century [Browder, 1965]. In this work, we investigate the variational inequality problem in the following form:

### v2

@@
Shuffling Heuristic in Variational Inequalities: Establishing New Convergence Guarantees Daniil Medyakov1111Corresponding author: medyakovd3@gmail.com, Gleb Molodtsov1, 2, Grigoriy Evseev1, Egor Petrov1, Aleksandr Beznosikov1, 2 1Moscow Center for Advanced Studies 2Innopolis University Variational inequalities have gained significant attention in machine learning and optimization research. While stochastic methods for solving these problems typically assume independent data sampling, we investigate an alternative approach – the shuffling heuristic. This strategy involves permuting the dataset before sequential processing, ensuring equal consideration of all data points. Despite its practical utility, theoretical guarantees for shuffling in variational inequalities remain unexplored. We address this gap by providing the first theoretical convergence estimates for shuffling methods in this context. Our analysis establishes rigorous bounds and convergence rates, extending the theoretical framework for this important class of algorithms. We validate our findings through extensive experiments on diverse benchmark variational inequality problems, demonstrating faster convergence of shuffling methods compared to independent sampling approaches.
@@ 1  Introduction
In addition to stochastic methods, various heuristics for selecting the $i_{t}$-th index at each iteration of algorithms exist. Careful analysis of these strategies may lead to the development of more robust and efficient algorithms. In this paper, we explore the shuffling heuristic [Mishchenko et al., 2020a; Safran and Shamir, 2020; Koloskova et al., 2023; Malinovsky et al., 2023]. Unlike the random and independent selection of the index $i_{t}$ at each iteration, which is common in classic stochastic methods, this heuristic adopts a more practical approach. Specifically, it involves permuting the sequence of indices $\{1,\ldots,n\}$, where $n$ is the number of data samples (4), and then selecting the index corresponding to the iteration number during the algorithm’s execution. This approach ensures that during one epoch of training, we take a step for each operator, and only once. There are several shuffling techniques available. Among the most popular are Random Reshuffling ($R$) [Gürbüzbalaban et al., 2021; Haochen and Sra, 2019; Nagaraj et al., 2019], where data is shuffled before each epoch; Shuffle Once (SO) [Safran and Shamir, 2020; Rajput et al., 2020], where shuffling occurs once before the start of training; and Cyclic permutation [Mangasarian and Solodov, 1993; Bertsekas and Tsitsiklis, 2000; Nedic and Bertsekas, 2001; Li et al., 2019], where data is accessed deterministically in a cyclic order.
@@
  • Columns: Sampling = Deterministic, if considered non-stochastic method, Independent, if method uses independent choice of operator’s indices, $R$ / SO if method uses shuffling heuristic, Assumption = assumption on operator $F$, VR? = whether the method uses variance reduction technique.
@@
In all these papers, the estimates were obtained through the formulation with an independent choice of the indices of the operator at each step of the algorithm. Regarding the shuffling heuristic, numerous studies explore methods suitable for addressing classic finite-sum minimization problems. In the work [Mishchenko et al., 2020a], the authors examined a classic SGD algorithm, and by introducing a new concept of variance specific to RR/SO, they matched the lower bounds in such scenarios. In the work [Malinovsky et al., 2023], the SVRG method with $R$ was considered. The authors actively utilized results from the work [Mishchenko et al., 2020a] and obtained improved rates. Additionally, there is a set of studies that considered methods incorporating the VR technique in the shuffling setup [Huang et al., 2021; Mokhtari et al., 2018; Ying et al., 2020]. However, there are currently no papers that employ the shuffling setting to solve variational inequalities. We aim to fill this gap.
@@ 3.2 Extragradient with Variance Reduction
Now we utilize the variance reduction technique, which enhances the convergence of the algorithms by diminishing the impact of random fluctuations. This approach was not employed in the previously presented algorithms. We introduce a variant of the RR/SO Extragradient with a variance reduction algorithm (Algorithm 3.2) and provide the convergence results for this method. In the work [Malinovsky et al., 2023], where shuffling is investigated in variance reduction methods, the authors utilize a more classic version and compute $F(\omega_{s}^{t})$ at the beginning of each epoch. We consider an alternative and compute this full operator randomly with probability pp. We set $p=\frac{1}{n}$, which implies that, on average, the full operator is updated once per epoch. Note that this choice does not increase the oracle complexity.
@@
We remove the variance that arose in Theorem 1 and obtain linear convergence. Even though we obtain worse estimates than those in works that also use the variance reduction technique, such as [Alacaoglu and Malitsky, 2022; Alacaoglu et al., 2021; Chavdarova et al., 2019; Palaniappan and Bach, 2016] (see Table 1), there is a distinct explanation for this. According to current theory, methods with the shuffling heuristic are inferior to methods with independent sampling for the variance reduction methods [Malinovsky et al., 2023]. In the prior works, where the minimization of the strongly-convex objective with shuffling was considered, the following rates are obtained: $\mathcal{\widetilde{O}}\left(n^{2}\frac{L^{2}}{\mu^{2}}\right)$ [Gurbuzbalaban et al., 2017], $\mathcal{\widetilde{O}}\left(n\frac{L^{2}}{\mu^{2}}\right)$ [Ying et al., 2020], $\mathcal{\widetilde{O}}\left(n\frac{L^{\nicefrac{{3}}{{2}}}}{\mu^{\nicefrac{{3}}{{2}}}}\right)$ [Malinovsky et al., 2023]. At the same time, in classic minimization, variance reduction methods provide $\mathcal{\widetilde{O}}\left(n\frac{L}{\mu}\right)$ rates [Gorbunov et al., 2020]. Thus, it remains an open question whether it is possible to obtain theoretical convergence estimates for methods using shuffling heuristics that are equivalent to those for methods with independent index selection for the VI problem. Additionally, in the course of the work, no theoretical differences are revealed in the SO and $R$ techniques concerning the problem (1).
@@ 4.1 Image Denoising
where $p$ is a dual variable, $\delta_{P}(p)$ is the indicator function of the set $P$ defined as: $P=\{p\in\mathcal{Y}:\|p(x)\|\leq 1\}.$ The indicator function $\delta_{P}(p)$ is defined as zero if $p$ belongs to the set $P$, and infinity otherwise. We define operator $\nabla u$ as the difference between neighboring pixels in the grid, both horizontally and vertically, normalized by the step of the grid $h$. This formulation represents a saddle point problem, where we seek to minimize the first term with respect to $u$ while simultaneously maximizing the second term with respect to pp. Using duality, we can write the final formulation of the problem as
@@
We configure the training with a fixed learning rate $\gamma=5\times 10^{-3}$ and a batch size of 88. The probability parameter $p=\frac{1}{n}$ is determined based on the effective dataset size. However, for computational efficiency, we adapt it to the mini-batch context. Training is conducted with random state 5050.
@@ 4.2 Adversarial Training
As shown in the plots, Extragradient and Extragradient with the VR algorithms using the independent choice of indices demonstrate worse performance compared to those using shuffling methods. In this series of experiments, $R$ exhibits better performance than other shuffling methods and significantly outperforms the non-shuffled versions.
Hyperparameter Details. We configure the training with a fixed learning rate $\gamma=0.01$ and a batch size of $4,4,16$ on mushrooms, a9a and w8a datasets, respectively. The probability parameter $p=\frac{1}{n}$ is determined based on the effective dataset size; however, for computational efficiency, we adapt it to the mini-batch context. Training is conducted with random state 5050.

### v3

@@
BRAIn Lab[Uncaptioned image] Shuffling Heuristic in Variational Inequalities: Establishing New Convergence Guarantees Daniil Medyakov1,2111Corresponding author: medyakovd3@gmail.com, Gleb Molodtsov1,2, Grigoriy Evseev1, Egor Petrov1, Aleksandr Beznosikov1,2 1Basic Research of Artificial Intelligence Laboratory (BRAIn Lab) 2Federated Learning Problems Laboratory Variational inequalities have gained significant attention in machine learning and optimization research. While stochastic methods for solving these problems typically assume independent data sampling, we investigate an alternative approach – the shuffling heuristic. This strategy involves permuting the dataset before sequential processing, ensuring equal consideration of all data points. Despite its practical utility, theoretical guarantees for shuffling in variational inequalities remain unexplored. We address this gap by providing the first theoretical convergence estimates for shuffling methods in this context. Our analysis establishes rigorous bounds and convergence rates, extending the theoretical framework for this important class of algorithms. We validate our findings through extensive experiments on diverse benchmark variational inequality problems, demonstrating faster convergence of shuffling methods compared to independent sampling approaches.
@@ 1  Introduction
In addition to stochastic methods, various heuristics for selecting the $i_{t}$-th index at each iteration of algorithms exist. Careful analysis of these strategies may lead to the development of more robust and efficient algorithms. In this paper, we explore the shuffling heuristic [Mishchenko et al., 2020a; Safran and Shamir, 2020; Koloskova et al., 2023; Malinovsky et al., 2023]. Unlike the random and independent selection of the index $i_{t}$ at each iteration, which is common in classic stochastic methods, this heuristic adopts a more practical approach. Specifically, it involves permuting the sequence of indices $\{1,\ldots,n\}$, where $n$ is the number of data samples (4), and then selecting the index corresponding to the iteration number during the algorithm’s execution. This approach ensures that during one epoch of training, we take a step for each operator, and only once. There are several shuffling techniques available. Among the most popular are Random Reshuffling (RR) [Gürbüzbalaban et al., 2021; Haochen and Sra, 2019; Nagaraj et al., 2019], where data is shuffled before each epoch; Shuffle Once (SO) [Safran and Shamir, 2020; Rajput et al., 2020], where shuffling occurs once before the start of training; and Cyclic permutation [Mangasarian and Solodov, 1993; Bertsekas and Tsitsiklis, 2000; Nedic and Bertsekas, 2001; Li et al., 2019], where data is accessed deterministically in a cyclic order.
@@
  • Columns: Sampling = Deterministic, if considered non-stochastic method, Independent, if method uses independent choice of operator’s indices, RR / SO if method uses shuffling heuristic, Assumption = assumption on operator $F$, VR? = whether the method uses variance reduction technique.
@@
In all these papers, the estimates were obtained through the formulation with an independent choice of the indices of the operator at each step of the algorithm. Regarding the shuffling heuristic, numerous studies explore methods suitable for addressing classic finite-sum minimization problems. In the work [Mishchenko et al., 2020a], the authors examined a classic SGD algorithm, and by introducing a new concept of variance specific to RR/SO, they matched the lower bounds in such scenarios. In the work [Malinovsky et al., 2023], the SVRG method with RR was considered. The authors actively utilized results from the work [Mishchenko et al., 2020a] and obtained improved rates. Additionally, there is a set of studies that considered methods incorporating the VR technique in the shuffling setup [Huang et al., 2021; Mokhtari et al., 2018; Ying et al., 2020]. However, there are currently no papers that employ the shuffling setting to solve variational inequalities. We aim to fill this gap.
@@ 3.2 Extragradient with Variance Reduction
Now we utilize the variance reduction technique, which enhances the convergence of the algorithms by diminishing the impact of random fluctuations. This approach was not employed in the previously presented algorithms. We introduce a variant of the RR/SO Extragradient with a variance reduction algorithm (Algorithm 3.2) and provide the convergence results for this method. In the work [Malinovsky et al., 2023], where shuffling is investigated in variance reduction methods, the authors utilize a more classic version and compute $F(\omega_{s}^{t})$ at the beginning of each epoch. We consider an alternative and compute this full operator randomly with probability $p$. We set $p=\frac{1}{n}$, which implies that, on average, the full operator is updated once per epoch. Note that this choice does not increase the oracle complexity.
@@
We remove the variance that arose in Theorem 1 and obtain linear convergence. Even though we obtain worse estimates than those in works that also use the variance reduction technique, such as [Alacaoglu and Malitsky, 2022; Alacaoglu et al., 2021; Chavdarova et al., 2019; Palaniappan and Bach, 2016] (see Table 1), there is a distinct explanation for this. According to current theory, methods with the shuffling heuristic are inferior to methods with independent sampling for the variance reduction methods [Malinovsky et al., 2023]. In the prior works, where the minimization of the strongly-convex objective with shuffling was considered, the following rates are obtained: $\mathcal{\widetilde{O}}\left(n^{2}\frac{L^{2}}{\mu^{2}}\right)$ [Gurbuzbalaban et al., 2017], $\mathcal{\widetilde{O}}\left(n\frac{L^{2}}{\mu^{2}}\right)$ [Ying et al., 2020], $\mathcal{\widetilde{O}}\left(n\frac{L^{\nicefrac{{3}}{{2}}}}{\mu^{\nicefrac{{3}}{{2}}}}\right)$ [Malinovsky et al., 2023]. At the same time, in classic minimization, variance reduction methods provide $\mathcal{\widetilde{O}}\left(n\frac{L}{\mu}\right)$ rates [Gorbunov et al., 2020]. Thus, it remains an open question whether it is possible to obtain theoretical convergence estimates for methods using shuffling heuristics that are equivalent to those for methods with independent index selection for the VI problem. Additionally, in the course of the work, no theoretical differences are revealed in the SO and RR techniques concerning the problem (1).
@@ 4.1 Image Denoising
where $p$ is a dual variable, $\delta_{P}(p)$ is the indicator function of the set $P$ defined as: $P=\{p\in\mathcal{Y}:\|p(x)\|\leq 1\}.$ The indicator function $\delta_{P}(p)$ is defined as zero if $p$ belongs to the set $P$, and infinity otherwise. We define operator $\nabla u$ as the difference between neighboring pixels in the grid, both horizontally and vertically, normalized by the step of the grid $h$. This formulation represents a saddle point problem, where we seek to minimize the first term with respect to $u$ while simultaneously maximizing the second term with respect to $p$. Using duality, we can write the final formulation of the problem as
@@
We configure the training with a fixed learning rate $\gamma=5\times 10^{-3}$ and a batch size of $8$. The probability parameter $p=\frac{1}{n}$ is determined based on the effective dataset size. However, for computational efficiency, we adapt it to the mini-batch context. Training is conducted with random state $50$.
@@ 4.2 Adversarial Training
As shown in the plots, Extragradient and Extragradient with the VR algorithms using the independent choice of indices demonstrate worse performance compared to those using shuffling methods. In this series of experiments, RR exhibits better performance than other shuffling methods and significantly outperforms the non-shuffled versions.
Hyperparameter Details. We configure the training with a fixed learning rate $\gamma=0.01$ and a batch size of $4,4,16$ on mushrooms, a9a and w8a datasets, respectively. The probability parameter $p=\frac{1}{n}$ is determined based on the effective dataset size; however, for computational efficiency, we adapt it to the mini-batch context. Training is conducted with random state $50$.